MsBrains (ms-brains):

Can somebody help check 12 of my practice questions? I need to make sure they are correct to prepare for class. I will fan and medal

MsBrains (ms-brains):

#8 I don't know how to solve these kind of problems, So my best guess is A.

Vocaloid (vocaloid):

well, first we pick an arbitrary number x then, we know that the ratio of 5x to 3x is 5 to 3, with me so far?

MsBrains (ms-brains):

Yeah

Vocaloid (vocaloid):

good, and the sum is 16 so 5x + 3x = 16 solve for x

MsBrains (ms-brains):

x=2?

Vocaloid (vocaloid):

good. now we work backwards and find 5x and 3x

MsBrains (ms-brains):

You lost me there, how do we work backwards?

Vocaloid (vocaloid):

well we just found out that x = 2 so what are 5x and 3x equal to?

MsBrains (ms-brains):

10 and 6?

Vocaloid (vocaloid):

yup so 10 and 6 are our answers

MsBrains (ms-brains):

Awesome, thanks so much! You are really helping me learn a bunch

Vocaloid (vocaloid):

uh you just posted the same question

MsBrains (ms-brains):

Oh! sorry give me second xD

OpenStudy (benlindquist):

lol

MsBrains (ms-brains):

#9. A?

OpenStudy (benlindquist):

Looks right, ?

Vocaloid (vocaloid):

hm not quite

OpenStudy (benlindquist):

do you add? divide?

Vocaloid (vocaloid):

find: 100*(new height-old height)/(old height)

MsBrains (ms-brains):

100*120/15 = 800

Vocaloid (vocaloid):

nope

Vocaloid (vocaloid):

new height - old height = ?

MsBrains (ms-brains):

Wait, so its not 100*120/15? It is 100*120-15?

Vocaloid (vocaloid):

neither

Vocaloid (vocaloid):

100*(new height-old height)/(old height) = 100*(120-15)/15 = ?

MsBrains (ms-brains):

100*(120-15)/15 = 700%

Vocaloid (vocaloid):

yes! [if you're wondering why we subtract new height - old height, it's asking for the INCREASE, so we need to know how much the height increased by first]

Vocaloid (vocaloid):

[this concept actually confuses a lot of people. "percent increase" is different than "percent of"]

MsBrains (ms-brains):

B?

Vocaloid (vocaloid):

not quite well, remember how we went from 3.1 to 310% by multiplying by 100? in this situation we're doing the opposite, so we divide by 100

MsBrains (ms-brains):

So A then? 0.0001

Vocaloid (vocaloid):

yup

MsBrains (ms-brains):

C?

Vocaloid (vocaloid):

nope (remember that "percent" means per 100, not 200 just divide 20/200

MsBrains (ms-brains):

So it;s 20%?

Vocaloid (vocaloid):

no, it's not 20%

Vocaloid (vocaloid):

what is 20 divided by 200?

MsBrains (ms-brains):

10%

Vocaloid (vocaloid):

good

MsBrains (ms-brains):

Last one. My best guess on this is A. I Don't know how to solve it. ;-;

Vocaloid (vocaloid):

simple interest means he earns 4% of 5500 each year so, first find 4% of 5500

MsBrains (ms-brains):

220%

Vocaloid (vocaloid):

just 220, no percent sign now multiply that by 5

MsBrains (ms-brains):

1100

MsBrains (ms-brains):

So it was A xD

MsBrains (ms-brains):

Right? o.O

Vocaloid (vocaloid):

yeah I didn't really look at the answer choices beforehand :P

MsBrains (ms-brains):

Thanks so much for the help. :)

MsBrains (ms-brains):

We're all done. :)